OpenStudy (anonymous):

what is the slope of the line described by the equation? 4x+6y=12

OpenStudy (imtiaz7):

we'll use slpe intercept form that is: y=mx+b from this we get: 6y=12-4x y=12/6 - 4x/6 y=2-2/3x From above equation we conclude: m=slope=-2/3